# Overcoming Reward Overoptimization via Adversarial Policy Optimization with Lightweight Uncertainty Estimation

### v1

@@ Abstract
We introduce Adversarial Policy Optimization (AdvPO), a novel solution to the pervasive issue of reward over-optimization in Reinforcement Learning from Human Feedback (RLHF) for Large Language Models (LLMs). Over-optimization occurs when a reward model serves as an imperfect proxy for human preference, and RL-driven policy optimization erroneously exploits reward inaccuracies. In this paper, we begin by introducing a lightweight way to quantify uncertainties in rewards, relying solely on the last layer embeddings of the reward model, without the need for computationally expensive reward ensembles. AdvPO then addresses a distributionally robust optimization problem centred around the confidence interval of the reward model’s predictions for policy improvement. Through comprehensive experiments on the Anthropic HH and TL;DR summarization datasets, we illustrate the efficacy of AdvPO in mitigating the overoptimization issue, consequently resulting in enhanced performance as evaluated through human-assisted evaluation.
@@ 1 Introduction
However, a critical issue arises in this process: the reward model, built from a finite dataset of human preferences, often fails to accurately represent the underlying human preference. This approximation error, worsened by the distribution shifts during policy updates [39], leads to unreliable rewards during the RL stage. This directly causes the phenomenon of ‘reward over-optimization’, wherein the LLM exploits erroneous high-reward states, artificially inflating the estimated proxy reward, while the ground-truth reward decreases [12, 9, 11]. This misalignment leads LLMs to prioritize reward maximization over actual content quality and user alignment, as seen in phenomena like ‘length bias’, where longer responses are erroneously favoured regardless of their relevance [33].
Current mitigation strategies against reward over-optimization, as proposed in [9, 11], focus on penalizing samples with high reward uncertainty during RL-based policy training. These approaches utilize ensembles of different reward models trained with different seeds, either during pre-training or fine-tuning phases, and leverage the variance in estimated rewards across ensembles to quantify uncertainty. Nevertheless, while in theory, these uncertainty estimates can identify unreliable rewards, the computational cost associated with training and maintaining multiple reward models in memory during policy optimization makes them impractical in real-world settings. What is worse, this computational constraint also hinders achieving the maximum performance potential, given that the “scaling laws” generally advocate for larger reward models [12, 41].
Our work builds upon the recent advancements in understanding the representation learning abilities in LLMs, particularly the last layer of a neural network. Numerous previous studies have shown that the last layer contains crucial information about the network’s prediction deviation from original training data [44, 17, 5, 31]. Inspired by these findings, we begin by inspecting lightweight uncertainty estimation methods in reward modelling, relying solely on the last layer embedding of the reward model. Subsequently, we propose Adversarial Policy Optimization (AdvPO), addressing overoptimization by solving a distributionally robust optimization problem centred around the reward model’s prediction confidence interval for policy improvement. This approach has proven to be a less conservative strategy for leveraging uncertainty in policy optimization compared to the utilization of uncertainty in previous work addressing overoptimization [9, 11]. In a variety of experiments, we show how AdvPO  effectively mitigates reward over-optimization without the computational burden as in the ensemble-based methods. Additionally, we demonstrate that AdvPO results in an improved policy overall, as evaluated through human-assisted assessments.
Our contributions are threefold:
@@
    Firstly, we conduct extensive experiments to demonstrate the effectiveness of last-layer embeddings, combined with off-the-shelf uncertainty estimation methods, in signalling over-optimization. We compare these against ensembles of comparable size and illustrate the advantage of our lightweight approach.
@@
    Secondly, we introduce Adversarial Policy Optimization algorithm (AdvPO) that adversarially searches around the reward model’s prediction confidence interval for policy optimization. We show that AdvPO is less conservative compared to previous work addressing over-optimization [9, 11].
@@ 2.1 Reinforcement Learning from Human Feedback
Before diving into our solution to remediate the reward overoptimization issue in Reinforcement Learning from Human Feedback (RLHF), we start by providing an overview of the most popularly studied RLHF workflow [25]. This helps us establish the notations and conceptual groundwork necessary for understanding our contributions. RLHF consists of three main steps: 1) Supervised Fine Tuning, 2) Reward Modelling, and 3) RL optimization.
Supervised Fine Tuning.
RLHF typically begins with Supervised Fine Tuning (SFT), which fine-tunes a pre-trained LLM through supervised learning on high-quality samples from downstream tasks, such as summarization or dialogue generation. We denote the resulting model as $\pi_{\rm SFT}$.
Reward Modelling.
The second phase of RLHF involves learning a reward model to capture human preferences through annotated preference data $D=\{(x^{i},y_{c}^{i},y_{r}^{i})\}_{i=1}^{N}$ where $y_{c}^{i}$ and $y_{r}^{i}$ denote the chosen and rejected responses to prompt $x^{i}$. The preferences are assumed to be generated by some unknown reward model $r^{*}(x,y)$ following the Bradley-Terry (BT) model [3]:
*$\mathbb{P}^{*}(y_{c}\succ y_{r}|x)=\frac{\exp(r^{*}(x,y_{c}))}{\exp(r^{*}(x,y_{c}))+\exp(r^{*}(x,y_{r}))}.$
Typically, a reward model $r_{\varphi}(x,y)$ is parameterized based on an LLM (usually $\pi_{\rm SFT}$), with an additional linear layer added to the final transformer layer that projects the last embedding layer to obtain a scalar reward. To be more specific, let $e(x,y):\mathcal{X}\times\mathcal{Y}\rightarrow\mathbb{R}^{d}$ denote the last layer embedding of the prompt and response pair $(x,y)$, and $\phi:\mathbb{R}^{d}\rightarrow\mathbb{R}$ denote the additional projection layer. We define the reward model as $r_{\varphi}(x,y):=\phi^{\top}e(x,y)$, where $\varphi$ includes all the tunable parameters in $\phi$ and $e(x,y)$.
Given the annotated preference data $D$, the reward model $r_{\varphi}$ is trained to assign higher rewards to the chosen answer $y_{c}$ compared to the rejected one $y_{r}$, by minimizing the negative log-likelihood under the BT model:
@@
RL optimization.
Lastly, the learned reward model $r_{\varphi}(x,y)$ is employed to provide feedback in an RL policy optimization phase (e.g., PPO [32]). Intuitively, the aim is to learn a policy $\pi_{\theta}$ that maximizes the reward $r_{\varphi}$ while not drifting too far away from $\pi_{\rm SFT}$:
$\textstyle\max_{\pi_{\theta}}\mathbb{E}_{x\sim D,y\sim\pi_{\theta}}\left[r_{\varphi}(x,y)\right]-\beta\mathbb{D}_{\text{KL}}\left[\pi_{\theta}(y|x)\middle\|\pi_{\rm SFT}(y|x)\right],$
@@ 2.2 Reward OverOptimization
A notable limitation of RLHF lies in the fact that the RL optimization process relies on the estimated reward $r_{\varphi}$, as opposed to the oracle/gold reward $r^{*}$. Though widely adopted, it overlooks the potential discrepancies between $r_{\varphi}$ and $r^{*}$, which may arise due to inaccuracies during the reward model training. Empirical studies have reported that the RL optimization tends to ‘hack’ the reward such that while the estimated reward (i.e., proxy reward) increases, the oracle/gold reward decreases. This phenomenon is referred to as overoptimization [12, 9, 11, 34, 26].
To mitigate this problem, in addition to the KL penalty in the original RL objective, several recent studies [9, 11] propose to leverage an ensemble of $K$ reward models $\{r_{\varphi_{k}}\}^{K}_{k=1}$. Given a prompt $x$ and its response $y$, these methods use the variance of rewards from different reward models to measure uncertainty in the estimated reward, i.e., $U_{x,y}={\rm Var}(\{r_{\varphi_{k}}(x,y)\}_{k=1}^{K})$. The reward is then penalized by the uncertainty before feeding into policy optimization:
@@
where $\gamma$ controls the degree of uncertainty-based penalty. Intuitively, samples with high uncertainty during policy training are penalized to reduce the risk of overoptimizing imperfect reward signals.
However, the use of reward ensembles, especially with LLMs, leads to significant computational overhead. This is due to the necessity of maintaining multiple reward models in memory throughout the policy optimization process, where each model is a separate LLM with billions of parameters. What is worse, it imposes limitations on the performance ceiling that can be achieved under given computational resources, as larger reward models are especially more powerful as suggested by the scaling law [12]. Thus, an intriguing question arises: Can we use lightweight uncertainty estimation in reward modelling to handle the overoptimization issue without burdensome ensembles?
@@ 3 Lightweight Uncertainty Estimation
In this section, we first argue that the final layer embedding of the reward model encapsulates crucial information regarding uncertainties in the model’s predictions. Following this, we propose a series of computationally efficient techniques for estimating the uncertainties associated with the predicted rewards, solely based on the final layer embeddings.
@@ 3.1 Connection between last layer embeddings and reward uncertainty
We begin by exploring a high-level intuition about how the final layer embeddings of a reward model can facilitate the estimation of uncertainties in the model’s output.
As discussed in Section 2.1, reward modelling can be decomposed into two parts: (1) learning a good representation $e(x,y)$ for the prompt and response pair through a pre-trained LLM; (2) projecting the learnt representation to yield a scalar reward using a mapping $\phi$. Very importantly, the training of LLMs on extensive text corpora, coupled with their vast number of parameters, enables these models to develop versatile representations that can even be used in zero/few-shot tasks [24, 41, 4], which demonstrate the adaptability of these representations.
However, the second part, which involves learning the projection weight $\phi$, is strictly tied to the preference data provided during the reward model training. Consequently, the reliability of rewards is closely linked to the accuracy and adaptability of the projection weight.
The above claim has been widely supported in previous work in deep neural networks [6, 17, 31, 44]. For instance, [17, 18, 19] demonstrate that by freezing the network up to its last layer and retraining only the projection head with a smaller data set, where spurious correlation is absent, it can greatly improve robustness of the neural network model against these spurious correlations. In the context of language models, recent experiments on weak-to-strong generalization [5] further reinforce this claim. Their findings reveal that, despite the fine-tuning of an LLM’s last layer embedding with noisy labels from weak supervision, the model can still excel in later classification tasks if the projection weight is accurately derived from ground-truth labels. This highlights the adaptability and the rich information encapsulated in the embeddings, accessible through simple projection weights.
Building upon the notion of generalized representations with specialized projection weights, we now shift our focus to the last layer’s ability for uncertainty estimation. The projection weight is strictly estimated based on the preference data encountered during reward model training. Therefore, when evaluating the prompt and response pairs during the RL stage, the pairs might deviate from what was observed during reward training (suggesting it undergo distribution shift [39]), hence rendering the predicted rewards unreliable as they might be outside the support of training set. To this end, we show in the next section how the last layer embedding of a reward model, based on preference data, can act as a feature map for an underlying kernel. This kernel then allows us to determine whether new prompt response pairs are similar to the ones seen during training. If not, the corresponding uncertainty should be higher, and therefore penalized during policy optimization.

### v2

@@ Abstract
Reinforcement Learning from Human Feedback (RLHF) has been pivotal in aligning Large Language Models with human values but often suffers from overoptimization due to its reliance on a proxy reward model. To mitigate this limitation, we first propose a lightweight uncertainty quantification method that assesses the reliability of the proxy reward using only the last layer embeddings of the reward model. Enabled by this efficient uncertainty quantification method, we formulate AdvPO, a distributionally robust optimization procedure to tackle the reward overoptimization problem in RLHF. Through extensive experiments on the Anthropic HH and TL;DR summarization datasets, we verify the effectiveness of AdvPO in mitigating the overoptimization problem, resulting in enhanced RLHF performance as evaluated through human-assisted evaluation.
@@ 1 Introduction
However, a critical issue arises in this process: the reward model, built from a finite dataset of human preferences, often fails to accurately represent the underlying human preferences. This approximation error, worsened by the distribution shifts during policy updates [38], leads to unreliable rewards during the RL stage. This directly causes the phenomenon of reward ‘overoptimization’, wherein the LLM exploits erroneous high-reward states, artificially inflating the estimated proxy reward, while the ground-truth reward decreases [12, 9, 11].
Current mitigation strategies against reward overoptimization, as proposed in [9, 11, 45], focus on penalizing samples with high reward uncertainty during RL-based policy training. Specifically, [9, 11] quantify reward uncertainty by training an ensemble of reward models with different seeds during either the pre-training or fine-tuning phases. They then measure the variance in estimated rewards across the ensemble to assess uncertainty in the reward prediction. However, RL policy optimization necessitates maintaining multiple reward models in memory, rendering it impractical for large models in real-world applications and limiting the potential to achieve maximum performance, especially since ‘scaling laws’ typically favour larger reward models [12, 40].
To reduce the memory footprint, recent works [45] instead suggest training multiple LoRA-based reward models with a diversity penalty for uncertainty quantification. However, even though LoRA ensembles reduce memory requirements by only having to train/store an adapter, they still incur high training costs, as well as computational overhead during policy optimization. More specifically, during the PPO stage, [45] requires querying each LoRA ensemble to compute the reward and uncertainty for every sample, which can easily become a serious computational bottleneck.
In this paper, we first introduce a lightweight method for quantifying reward uncertainty in the RLHF pipeline, using only the last layer embeddings of the reward model. This approach is easily integrated into any existing trained reward models, making it generally applicable. Building on these uncertainty estimates about reward prediction, we then propose Adversarial Policy Optimization, AdvPO, a distributionally robust optimization procedure to counter overoptimization during policy improvement. AdvPO contrasts with previous sample-wise uncertainty penalization methods [9, 11, 45], for which we theoretically prove that AdvPO handles reward uncertainty in a less pessimistic manner. As a result, AdvPO is more effective at improving policy and mitigating overoptimization, which we empirically confirm in extensive experiments. These favourable results are further supported through human-assisted assessments.
To summarize, our contributions are threefold:
@@
    Firstly, we introduce a lightweight method to quantify reward uncertainty using only the last layer embeddings of the reward model into the RLHF pipeline. Extensive experiments confirm its effectiveness in identifying reward uncertainties and signalling overoptimization.
@@
    Secondly, we introduce the Adversarial Policy Optimization (AdvPO), built on our efficient uncertainty estimates, to adversarially target the reward model’s prediction confidence interval for policy optimization. AdvPO is proven to be less pessimistic than sample-wise uncertainty penalization methods [9, 11], thus more effective at enhancing the policy and mitigating overoptimization.
@@ 2.1 Reinforcement Learning from Human Feedback
We start by providing an overview of the RLHF workflow [26]. This helps us establish the notations and conceptual groundwork necessary for understanding our contributions. RLHF consists of three main steps: 1) Supervised Fine Tuning, 2) Reward Modelling, and 3) RL optimization.
Supervised Fine Tuning. RLHF typically begins with Supervised Fine Tuning (SFT), which fine-tunes a pre-trained LLM through supervised learning on high-quality samples from downstream tasks, such as summarization or dialogue generation. We denote the resulting model as $\pi_{\rm SFT}$.
Reward Modelling. The second phase of RLHF involves learning a reward model to capture human preferences through annotated data $D=\{(x^{i},y_{c}^{i},y_{r}^{i})\}_{i=1}^{N}$ where $y_{c}^{i}$ and $y_{r}^{i}$ denote the chosen and rejected responses to prompt $x^{i}$. The preferences are assumed to be generated by some unknown reward model $r^{*}(x,y)$ following the Bradley-Terry (BT) model [3]:
$\mathbb{P}^{*}(y_{c}\succ y_{r}|x)=\frac{\exp(r^{*}(x,y_{c}))}{\exp(r^{*}(x,y_{c}))+\exp(r^{*}(x,y_{r}))}.$
Typically, a reward model $r_{\varphi}(x,y)$ is initialized from a pretrained LLM (usually $\pi_{\rm SFT}$), with an additional projection layer added to map the last embedding layer to a scalar reward. To be more specific, let $e(x,y):\mathcal{X}\times\mathcal{Y}\rightarrow\mathbb{R}^{d}$ denote the last layer embedding of the prompt and response pair $(x,y)$, and $\phi:\mathbb{R}^{d}\rightarrow\mathbb{R}$ denote the additional projection layer. We define the reward model as $r_{\varphi}(x,y):=\phi^{\top}e(x,y)$, where $\varphi$ includes all the tunable parameters in $\phi$ and $e(x,y)$.
Given the annotated preference data $D$, the reward model $r_{\varphi}$ is trained to assign higher reward to the chosen response $y_{c}$ compared to the rejected one $y_{r}$, by minimizing the negative log-likelihood under the BT model:
@@
RL optimization. Lastly, the learned reward model $r_{\varphi}(x,y)$ is employed to guide the RL policy optimization phase (e.g., PPO [32]). Intuitively, the aim is to learn a policy $\pi_{\theta}$ that maximizes the reward $r_{\varphi}$ while not drifting too far away from $\pi_{\rm SFT}$:
$\textstyle\max_{\pi_{\theta}}\mathbb{E}_{x\sim D,y\sim\pi_{\theta}}\left[r_{\varphi}(x,y)\right]-\beta\mathbb{D}_{\text{KL}}\left[\pi_{\theta}(y|x)\middle\|\pi_{\rm SFT}(y|x)\right],$ (2)
@@ 2.2 Reward Overoptimization
A notable limitation of RLHF lies in the fact that the RL process relies on the estimated reward $r_{\varphi}$, as opposed to the oracle/gold reward $r^{*}$. Though widely adopted, it overlooks the potential discrepancies between $r_{\varphi}$ and $r^{*}$, which may arise due to inaccuracies during the reward model training. Empirical studies have reported that the RL stage tends to ‘hack’ the reward such that while the estimated reward (i.e., proxy reward) increases, the oracle/gold reward decreases. This phenomenon is referred to as overoptimization [12, 9, 11, 33, 27].
To mitigate this problem, in addition to the KL penalty in the original RL objective, several recent studies [9, 11, 45] propose to leverage an ensemble of $K$ reward models $\{r_{\varphi_{k}}\}^{K}_{k=1}$. Given a prompt $x$ and its response $y$, these methods use the variance of rewards across different reward models to measure uncertainty in the estimated reward, i.e., $U_{x,y}={\rm Var}(\{r_{\varphi_{k}}(x,y)\}_{k=1}^{K})$. The reward is then penalized based on the sample-wise uncertainty before being used in policy optimization:
@@
where $\gamma$ controls the degree of uncertainty-based penalty. Intuitively, samples with high uncertainty during policy training are penalized to reduce the risk of being misled by imperfect rewards. However, as previously mentioned, reward ensembles that are trained either by fine-tuning entire LLMs [9, 11] or by using LoRA [45] incur additional memory and computational overhead. This is due to the need of maintaining multiple reward models in memory for policy learning.
Thus, an intriguing question arises: Can we quantify reward uncertainty in a lightweight manner that can be easily integrated into any trained reward models, thereby addressing the overoptimization issue without the need for burdensome ensembles? And in the following section, we provide a positive answer to this important question.
@@ 3 Lightweight Uncertainty Estimation
In this section, we introduce a lightweight uncertainty quantification method, based solely on the final layer embeddings of a reward model. We start by offering a high-level intuition on why the last layer embedding captures essential information about uncertainties in the reward model’s predictions. Following this, we present our lightweight uncertainty quantification method.
@@ 3.1 Connection between last layer embeddings and reward uncertainty
As discussed in Section 2.1, reward modelling can be decomposed into two parts: (1) learning a good representation $e(x,y)$ for the prompt and response pair through a pre-trained LLM; (2) projecting the learnt representation to a scalar reward using a mapping $\phi$. Very importantly, the training of LLMs on extensive text corpora, coupled with their vast number of parameters, enables these models to develop versatile representations that can even be used in zero/few-shot tasks [25, 40, 4], which demonstrate the generalizability of these representations.
However, the second part, which involves learning the projection weight $\phi$, is strictly tied to the preference data provided during the reward model training. Consequently, the reliability of predicted rewards is closely linked to the accuracy and generalizability of the projection weight.
The above claim has been widely supported in the deep learning literature [6, 17, 31, 43]. For instance, [17, 18, 19] demonstrate that by freezing the network up to its last layer and retraining only the projection head with a smaller data set, where spurious correlation is absent, it can greatly improve robustness of the neural network model against these spurious correlations. In the context of language models, recent experiments on weak-to-strong generalization [5] further reinforce this claim. Their findings reveal that even when fine-tuning an LLM’s last layer embedding with noisy labels from weak supervision, the model can still excel in subsequent classification tasks if the projection weight is accurately derived from ground-truth labels. This highlights the generalizability and the rich information encapsulated in the last layer representation, accessible by simple projection weights.
Building upon the notion of generalized representations with specialized projection weights, we now zoom our attention to the last layer’s ability to quantify the uncertainty of its output. The projection weight is strictly estimated based on the preference data encountered during reward model training. Therefore, when evaluating the prompt and response pairs during the RL stage, the pairs might deviate from what was observed during reward model training (suggesting a distribution shift [38]), hence rendering the predicted rewards unreliable.
In the next section, we show how the last layer embedding of a reward model, based on preference data, can act as a feature map for an underlying kernel (similarity measure). This kernel then allows us to determine whether new prompt response pairs are similar to the ones seen during training. If not, the corresponding uncertainty should be higher and penalized during policy optimization.